OpenStudy (anonymous):

how do you find the integral of 1/sqrt(9-x^2)

OpenStudy (anonymous):

u have formula for dis look into ur text book

OpenStudy (anonymous):

dx/sqre root(a^2-x^2)

OpenStudy (amistre64):

consider x = 3 sin(t) dx = 3 cos(t) dt

OpenStudy (amistre64):

\[\int\frac{1}{\sqrt{9-x^2}}dx\] \[x=3~sin(t)~:~sin^{-1}(\frac x3)=t\\ dx=3~cos(t)~dt\] \[\int\frac{3~cos(t)}{\sqrt{9-(3~sin(t))^2}}dt\] \[\int\frac{3~cos(t)}{\sqrt{9-9~sin^2(t)}}dt\] \[\int\frac{3~cos(t)}{3\sqrt{1-sin^2(t)}}dt\] \[\int\frac{3~cos(t)}{3~cos(t)}dt\] \[\int~dt=t+C\] \[sin^{-1}(\frac x3)+C\]